Steve traveled 600 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 1 hour less,. Find the speed of his vehicle
Answer by richwmiller(17219)

r*t=600,
(r+20)*(t-1)=600
r = 100 mph speed , t = 6 hours
r+20=120 t-1=5